OpenStudy (anonymous):

Please help, I'll give you a medal<3 The sample space for a roll of two number cubes in shown in the table (1,1), (1,2), (1,3), (1,4), (1/5), (1/6) (2,1), (2,2), (2,3), (2,4), (2,5), (2,6) (3,1), (3,2), (3,3), (3,4), (3,5), (3,6) (4,1), (4,2), (4,3), (4,4), (4,5), (4,6) (5,1), (5,2), (5,3), (5,4), (5,5), (5,6) (6,1), (6,2), (6,3), (6,4), (6,5), (6,6) The two numbers rolled can be added to get a sum. Find P(Sum is great than 5) A. 5/6 B. 13/18 C. 5/18 D. 1/3

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

count how many are greater than 5

OpenStudy (anonymous):

there are a lot, too many for me to write them all but you can count them then divide that number by \(36\)

OpenStudy (anonymous):

i only count 10 less than or equal to 5, leaving the rest 26 greater than 5 that makes it \(\frac{26}{36}\) which you can reduce

OpenStudy (anonymous):

is it 1/3?

OpenStudy (anonymous):

Wait..

OpenStudy (anonymous):

It's B. right? @satellite73

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

Crap... Lmao

OpenStudy (anonymous):

yes, it is B

OpenStudy (anonymous):

\[\frac{26}{36}=\frac{13}{18}\]

OpenStudy (anonymous):

Oh okay, good! Thank you! I'm finally getting how to do this!